1)   A glass window is exactly 20 cm by 30 cm at 10 0C. By how much would its area increase in cm2 when

 

its temperature is 40 0C? Take the linear expansivity of glass to = 9 X 10-6 / C0.


1
Expert's answer
2022-02-14T13:56:17-0500

Solution

Area A=20*30cm2 =0.06m2

T=10°C

A'=?

T=40°C

linear expansivity of glass to = 9 X 10-6/°C

So area becomes as

"A'=A(1+\alpha \Delta T) \\
=0.06(1+9*10^{-6}\times30) \\=0.06*1.00093\\=0.0600558m^2"